In an atom of hydrogen,the atom is exposed to a radiation of wavelenght 970.58Å and atom got excited.Find out the maximum number of waves can be obtained during de excitation.

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Hiiii 
The maximum no of waves excited wil be equal to ???
we know that 
1/Lambda = Rh (1/nf2 – 1/ni2
Rh = 1.0973 * 10-7
1/(970.58 * 10-10) * 1.0973 * 10-7   =  (1/nf2 – 1/ni2
There will be varable values of nf and ni